OpenStudy (anonymous):

find the exact value of sin tetha/2, Cos tetha = -4/5, tetha is in qudrant 3

OpenStudy (dumbcow):

use half angle identity \[\sin (\theta/2) = \sqrt{\frac{1-\cos \theta}{2}}\] since theta is in quad 3, sin(theta/2) will be positive

OpenStudy (anonymous):

But how do you find the value?

OpenStudy (dumbcow):

plug in value for cos(theta)